OpenStudy (anonymous):

I need help with 1 more question: Write the slope-intercept form of the equation of line passing through the point (4,6) and parallel to the line y= -6x-1

OpenStudy (anonymous):

parallel lines means same slope so the slope for our line will be -6 as well now line equation : y = y1 + m(x-x1) y1 = 6 x1 = 4 m = -6 y=6 + (-6)*(x-4) y =6 -6x + 24 y = -6x + 30

OpenStudy (anonymous):

thank you !

OpenStudy (anonymous):

sup abdu